OpenStudy (anonymous):

Given that the volume of water a pipe 500 m long with a volume of 98.187m^3, is 98187 liters. if water flows through the pipe at 0.8m/s, how many liters will flow through the pipe in 1 hour?? pls show working if possible, thanks :)

OpenStudy (anonymous):

The tricky part is that 0.8m/s is telling you the velocity at which water is coming out, to know how much water is coming out in Volume you need to determine the volumetric flow rate (units m^3/time). The missing piece of information is the area of the pipe. Liters in one hour = VolumetricFlow * time VolumetricFlow = PipeArea * 0.8 m/s * 3600 s/h = time = 1 hour Assuming the pipe is a cylinder PipeVolume = PipeArea * length PipeArea = PipeVolume/length = 98.187 m^3/500m = 0.1964 m^2 Substitution the pipe area in the equation for VolumetricFlow Liters in one hour = 565.63 m^3 The attached file has details of the calculation.

OpenStudy (anonymous):

oh wow touchy maths how did you do that

OpenStudy (anonymous):

but i didn't give you the entire info. the diameter is 0.5m :P

OpenStudy (anonymous):

you can calculate that no need for that

OpenStudy (anonymous):

is there any simpler wayy??

OpenStudy (anonymous):

I say find the cross section are pir^2

OpenStudy (anonymous):

that is 0.196375.

OpenStudy (anonymous):

then calculate time to travel along the pipe with the water's speed

OpenStudy (anonymous):

end to end

OpenStudy (anonymous):

ah calculate I don't like calculations I need your support

OpenStudy (anonymous):

500 m is distance and speed is .8m/s

OpenStudy (anonymous):

ah no need for calculations time is 625 seconds

OpenStudy (anonymous):

now in 625 second a volume of 98.something will flow through the pipe

OpenStudy (anonymous):

98187 litres

OpenStudy (anonymous):

I find it easier to work these physics problems if you start by defining the solution using words and work your way back. The solution is: Volume = Flow * time of flow time = 1 hour = 60 minutes = 3600 seconds ...

OpenStudy (anonymous):

ah there are 3600 seconds in an hour

OpenStudy (anonymous):

oh, wait, i think i got it :) 1liter can flow in 0.174 hours, so x liters will flow in 1 hour, and cross multiply. btw, touchymath, that, was, AMAZING!!

OpenStudy (anonymous):

Glad I could help! : )

OpenStudy (anonymous):

and the volume would be 5.76 l, right?? yay !!!

OpenStudy (anonymous):

ah no I think you got wrong somewhere

OpenStudy (anonymous):

Actually, there are 1000 liters in one m^3, so the answer is 565.63 (meters^3) = 565 630 liters

OpenStudy (anonymous):

3600 ------ x 98187 = 565557.12 litres 625

OpenStudy (anonymous):

ah that is what I get well both are equivalent

OpenStudy (anonymous):

anyways, whats the formula for finding the length of a line on a graph?

OpenStudy (anonymous):

\[d = \sqrt{(x_2-x_1)^2 +(y_2-y_1)^2}\]

OpenStudy (anonymous):

aha!! i was right!!! thanks a LOT!!!! :D